The population of a city t years from now is given by P=250,000e^0.04t,
Find the rate of change of population with respect to time t three years from now. (Round your answer to the nearest integer.)
(b)What is the actual change in population in the 4th year. (Round your answer to the nearest integer.)

Answer by Boreal(15235)

derivative of 250000e^0.04t=0.04(250000)*e^0.04t, evaluated at t=3
e^0.12=1.128, and without rounding, multiply by 0.04*250000=11,274.96, or 11275 people per year.
between 3 and 4 years, the change will be 250000(e^0.16-e^0.12)=11,503.50 or 11,504 people.